# Bridging the Gap Between Saliency Prediction and Image Quality Assessment thanks: This study was supported by Russian Science Foundation under grant 24-21-00172, https://rscf.ru/en/project/24-21-00172/

### v1

@@ V Conclusion
This study examined the relationship between saliency prediction and image-quality assessment. We propose a technique for extraction saliency maps from IQA models and empirically demonstrate their reliance on salient regions. Additionally, we propose a parameter-free dual-task learning approach without sacrificing quality. Finally, we curated a dataset of non-uniformly compressed images and performed large-scale comparison, which reveals that NR metrics are better suited for saliency-aware compression domain. We conclude that saliency can remarkably improve NSS-based methods, while learning-based methods face slight improvement.

### v2

@@ V Conclusion
This study examined the relationship between saliency prediction and image-quality assessment. We propose a technique for the extraction of saliency maps from IQA models and empirically demonstrate their reliance on salient regions. Additionally, we propose a parameter-free dual-task learning approach without sacrificing quality. Finally, we curated a dataset of non-uniformly compressed images and performed a large-scale comparison. We conclude that saliency can remarkably improve NSS-based methods, while learning-based methods face slight improvement since they already take saliency priors into account during conventional training.
Acknowledgments
This research was supported by Russian Science Foundation under grant 24-21-00172, https://rscf.ru/en/project/24-21-00172/ and was carried out using the MSU-270 supercomputer of Lomonosov Moscow State University.